# Towards Human-centered Explainable AI: A Survey of User Studies for Model Explanations

### v4

@@ 1 Introduction
[…]
.98! Trust [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] Understanding subjective [32, 33, 34, 35, 22, 36, 37, 12, 13, 17, 28, 16, 38, 39, 7, 40, 41, 42, 43, 14, 44, 31] objective [45, 40, 46, 32, 47, 48, 49, 50, 51, 52, 13, 12, 53, 54, 55, 56, 35, 39, 57, 22, 58, 59, 43, 60, 44] explanation model [61, 62, 63, 64, 46, 21, 65, 49] Usability workload [48, 21, 16, 3, 66] helpfulness [46, 48, 56, 65, 67, 13, 68, 45] satisfaction [47, 16, 7, 18, 6, 69, 1, 29, 19, 70] undesired behavior detection [53, 24, 71, 57, 72, 38, 73, 74, 75, 27, 76, 2, 77, 78, 79] ease of use and others [32, 71, 21, 13, 48, 80, 37, 81, 30, 65, 24, 3, 82, 83, 1, 66, 84, 85, 20, 86, 87] Human-AI Collaboration Performance [88, 10, 89, 25, 90, 91, 92, 13, 15, 93, 56, 30, 94, 39, 25, 53, 29, 95, 96, 43, 97]
@@
[…]
! Explainable AI User Study Keywords XAI, explainable AI, explanation, explainable, explanatory, interpretable, intelligible, black-box, machine learning, explainability, interpretability, intelligibility, explain attribution, feature user study, participant , human subject, empirical study, lab study, user evaluation, human evaluation
@@ 3.2 Foundations of User Studies
Based on a data-driven bibliometric analysis of the references in core papers, we highlight significant research topics within the "Foundational Domain" in Figure 1. It is evident that model explanations and interpretability are pivotal components. This includes papers that introduce explanation methods such as LIME [147], SHAP [148], and other attribution methods. These are a frequent subject of study in works measuring understanding and usability. Additionally, convolutional networks, which are commonly employed in experiments, use tools like GradCAM [149] and various saliency maps to generate model explanations. Notably, many research papers appear within the domain of recommender systems, because many XAI user studies are conducted in the context of recommendation solutions. The EU’s General Data Protection Regulation (GDPR) [150] is frequently mentioned in core papers due to the ongoing debate on the right to explanation” [151]. This debate has significantly influenced the shift in modern AI systems towards explainability. While the ultimate consumers of model explanations are humans, well-established research domains that focus on human understanding are underrepresented. For instance, only a few papers related to “Cognition” are cited compared to those on other algorithmic topics. Millecamp et al. [18] suggest enhancing XAI theory with insights from social sciences, including cognitive science and psychology. Given the scant references to psychology, it appears that only a handful of XAI user studies delve into evaluating XAI from a psychological standpoint. We highlight a nascent research domain of XAI frameworks based on human cognition and behavior theories [142]. This theoretical guidance can also offer conceptual tools for better evaluating XAI from user perspectives. More details about common references can be found in Appendix LABEL:app-appendix:foundation.
@@ 3.3 Impact of User Studies
Figure 1 presents applications that make use (and thus are the consumers) of the findings from core papers. We noticed that studies on user understanding and trust span a wide range of applications. For example, trust is frequently addressed in the contexts of medical diagnosis and transportation, indicating its significance in high-risk scenarios. Recommendation systems emerge as a primary focus in follow-up works. Papers on usability have a significant impact on fields like data visualization, software development, and education. In these areas, models frequently serve as tools to ease the burden on end users. Human-AI collaboration measures particularly promote the further development of robotics and or natural language processing. The prominence of recommendation systems in both foundational works and their impact implies that XAI is an integral component of contemporary recommendation systems. A comprehensive overview of the fundamental works and application domains can be found in Appendix LABEL:app-appendix:combined.
@@ 4.1 Models and Explanations
[…]
! White-box Black-box Other Feature- based local [21, 154, 48] [22, 12, 39] [6, 50] [21, 49, 55, 45] [34, 29, 92, 72] [39, 35, 40] [47, 65, 42] [54, 58, 57] [71, 56, 50] [40, 41, 89] [90, 25, 59] [95, 43, 60] global [12, 53, 74] [21, 50] [50] Example-based [12, 21, 43] [74, 6, 96] [52, 17, 57] [13, 40, 25] [32] (generative models) Counterfactual [12, 37] [82, 21] [27, 100] [57, 65] Concept-based [61, 62, 71] [67, 63, 64] [57, 100] Other [11, 88] [7, 10] [155, 1, 9, 15] [3, 13, 51] [3, 56, 58] [49, 55, 36] [28, 85, 16] [68, 33, 38]* [76, 8, 23] * [69, 70]* [2] [18, 20, 19] $\dagger$ [26, 84, 20] $\dagger$ [83, 66, 81] $\dagger$ [85, 30, 14] $\dagger$ [91, 5] $\dagger$
TABLE III: Models and explanations in core papers. Papers are categorized according to types of explanations (column) and types of models (row). *** denotes papers using recommendation systems as models; $\dagger$ denotes papers proposing novel interpretable interfaces as studied models.
Besides these four main types of explanations, there are other explanations such as rules [11, 88] or game strategies [7, 10] when AI plays games. More details about concrete models and explanations can be found in Appendix LABEL:app-sec:modelexplanations.
@@ 4.2.1 Trust
Trust measures used in much of the existing research can be divided into two groups: self-reported and observed trust [156]. Self-reported trust is commonly measured by asking users to fill out questionnaires whereas observed trust is quantified by humans’ agreement with the model’s decisions. In LABEL:app-tab:trust_measures in Appendix, trust measures in these two groups are listed. The agreement rate of users with the model decisions is commonly used [9, 12, 11, 25] as a measure of observed trust. Parallel to observed trust measurement, van der Waa et al. [157] ascribe the user’s alignment behaviors to the persuasive power of model explanations, i.e., the capacity to convince users to follow model decisions despite the correctness. As an extension, trust calibration is defined based on this measure. For example, a high agreement rate to wrongly made decisions represents overtrust, while a low agreement rate to correct decisions means undertrust [12]. In self-reported measurements, researchers either utilize well-developed questionnaires or self-designed ones, with the exception of [4] which conducts a semi-structured interview to explore user opinions. Several works [6, 13, 27, 11, 16, 17, 18, 19, 24] propose their own questionnaires. Among these, a subgroup [16, 13, 18, 19, 24] simply asks users to rate a single statement such as “I trust the system’s recommendation/decision”, which is named as one-dimensional trust by [8]. When deploying previously proposed questionnaires [2, 3, 21, 7, 23, 5, 8, 10, 22, 158], Trust in Automation [159] is the most commonly used one, in which the underlying constructs of trust between human and computerized systems are explored.
@@ 4.2.2 Understanding
An important goal of explanation techniques is to foster users’ understanding of complex ML systems. An important separation has to be made between users’ perceived understanding and their actual comprehension of the underlying model, as the two often do not agree [35, 40]. Cheng et al. [22] explicitly differentiate between objective understanding and self-reported understanding, which we term subjective understanding in this work. While subjective understanding is usually measured through questionnaires, measuring objective understanding requires a proxy task where the users’ understanding is put to a test. Additionally, user studies can be run to assess how well users can understand the explanation itself (and not the underlying model). This can be an important sanity check and is particularly used in the domain of conceptual explanations [160, 62], where the intelligibility of concepts needs to be verified. We refer to the third category as understanding of explanations but defer its detailed findings to Appendix LABEL:app-sec:understandingofexplanations.
Objective Understanding. Works in the subdomain of objective understanding deploy proxy tasks to verify users’ understanding of a model’s inner workings. The most commonly considered domain in works on understanding is finance [53, 48, 49, 40, 47, 35, 39] followed by image classification [13, 21, 52]. One of the most critical design choices when assessing objective understanding is the selection of a suitable proxy task. Doshi-Velez and Kim [120] argue that the task should “maintain the essence of the target application” that is anticipated. One of the most prominent tasks is forward simulation [141, 120]. This task demands subjects that are given an input to simulate, i.e., predict, the model’s output. The extent to which participants can successfully provide the model’s output is also referred to as simulatability [141]. However, scholars have designed many more tasks to quantify understanding and applied them across a variety of data modalities (cf. Table LABEL:app-tab:proxytasks in Appendix for an exhaustive listing).
@@ 4.2.3 Usability
Using model explanations to audit models is one purpose of explainability [130]. Some of the surveyed works study how model explanations can assist users in detecting undesired behaviors of models. These issues mainly include (perceived) unfairness in the model decision-making [38, 74, 78, 79], biases in models [72] or features [57], and wrong decisions (failures) [24] in the studied papers. A detailed summary of types of undesired behaviors is listed in Table VI. In the undesired behavior detection, the effectiveness of explanations is evaluated by objective performance measures, such as the number of bugs identified [71], the share of participants that identify a certain bias [57, First Experiment] or by the deviations between model predictions and human predictions for unusual samples [53]. The perception of users regarding fair treatment by a system has primarily been researched in high-stakes applications such as granting loans [27] or granting bail for criminal offenders [74, 75, 73]. For example, [74, 75, 73] investigate the fairness of COMPAS, a commercial criminal risk estimation tool that was used in the US to help make judicial bail decisions. It is also considered in everyday use-cases such as news [38] and music [77] recommendations, or possible career suggestions [76], where a bias in the underlying system can be to the detriment of the user. As the assessment of fairness is a very subjective matter, questions regarding perceived fairness are prevalent, e.g., “how the software made the prediction was fair” [74], which can be answered on 5- or 7-point Likert scales [74, 75, 27, 38, 2, 73]. Among these works, an effective explanation is the one that can either increase or decrease the fairness perceptions, since the aim of explanations is to show fairness or unfairness. An exhaustive overview of measures for usability is given in LABEL:app-tab:usability_measures of the Appendix.
@@ 4.3 Experimental Design and Analysis
[…]
! Experimental Design Between-Subjects Within-Subjects Mixed Papers [5, 7, 8, 12, 27, 59, 15] [17, 21, 22, 23, 25, 72] [28, 11, 40, 46, 32, 95] [53, 47, 49, 50, 51, 84] [56, 39, 38, 37, 36, 43] [90, 54, 30, 29, 92, 96] [57, 58, 12, 73, 38] [75, 76, 77, 78, 82] [1, 3, 4, 19, 9] [18, 70, 10, 24, 21] [26, 13, 52, 35, 91] [57, 78, 71, 81, 93] [69, 6, 67, 63, 62] [68, 41, 60, 14, 45] [2, 66, 13, 52, 16] [74, 28, 34, 64, 10] [33, 147, 83, 65, 12] [61]
@@ 5 Findings of User studies
In this section, we summarize the primary findings from the core papers. Table V lists findings with respect to four measured quantities. To build an overview of the findings, we divide papers according to their evaluation dimensions, i.e., the independent variables in the user studies. When using the presence of explanations as the evaluation aspect, the findings are summarized in Table V. The listed impacts using explanations are to be seen in comparison with a control group without explanations. Effects are divided into two groups: (1) Positive effects, for example, increasing user trust or understanding; (2) Non-positive effects: the effect can be negative, or not significantly positive (neural), or a mixture of different effects (e.g., feature-based explanations have positive effects but counterfactual explanations do not). Beyond the explanations themselves, other possible evaluation dimensions such as that might have an impact on the perception of XAI, for instance, AI technology literacy, model performance, or the dimensionality of the data. Instead of using the mere presence of explanations, many works compare different explanation techniques with each other (see Appendix LABEL:app-sec:appfindings for more details).
@@
As various research questions and findings are addressed in 97 core papers, many papers compare explanation types in order to choose a preferable one, it is not possible to cover all results in one table. Based on them, we outline some interesting trends in the effectiveness of explanations on user experience: (1) Explanations are effective in improving users’ subjective understanding; (2) The effectiveness of explanations in increasing user trust and usability of models is not clear; (3) Explanations are not good at convincing users that models are fair; (4) Interactivity of the model has positive impact on user trust, understanding and model usability. The first three statements can validated through the number of papers obtaining positive or non-positive effects in each category, while the last finding is extracted from LABEL:app-tab:findings2 in the Appendix, which details findings with on other independent variables. We encourage the reader to consider the short summary of primary findings in the tables and check for further details according to their specific interests. In the following section, we highlight some findings for each category of measurement.
[…]
! Evaluation Dimension: Explanations Effect of explanations compared to no explanations Positive Non-positive / Mixed Trust [13]: example-based, rule-based explanations [16]: example-based explanations for recommendations [27]: feature importance [10]: decision-tree explanation for policy [28]: explanation corpus given by researchers [25]: feature-based (saliency map), example-based explanations [8]: explanations for recommendations [6]: rationale-based, example-based and feature-based (best) explanations for online symptom checkers [15]: confidence scores [3]: positive in simulation but no improvement in real-word [1]: explanations for medical suggestions (Doctor XAI [155]) pos.for observed trust but insignificant for reported trust [12]: feature-based explanations increase appropriate trust slightly but counterfactual explanations inconclusively [22, 21, 24]: feature-based explanation, insignificant [11]: rule-based explanation, insignificant [15]: Shapley values, insignificant [29]: feature-based explanation, negative Understanding Obj. [53, 22] white-box model [40] feature importance, LIME (tabular) [46] counterfactuals+cues (audio) [50] manipulatability improved by white-box log. reg. [54] saliency maps (image) [59] saliency maps for bias detection and strategy identification [12] counterfactuals+feature importance [39]: SHAP, negative for black-box model (education domain) [39]: Insignificant difference btw. black-box and white-box models [40]: Prototypes, Anchors, LIME on textual data insignificant [46]: Counterfactuals and Concepts insignificant (audio data) [50]: Simulatability results insignificant for LIME, IG, surrogate model on BERT and Logistic Regression Model, Manipulatability insignificant for BERT [58]: Insignificant results for GRAD-CAM, Saliency Map, uncertainty scores in VQA [59] saliency maps for failure prediction (image) [60]: saliency maps, negative for a mix of three interpretation techniques in simulation task Sub. [13]: example-based, rule-based explanations [28]: explanation corpus given by researchers [12]: feature-, example- and counterfactual-based [38]: explanations provided by [167] for Facebook News Feed [16]: example- and feature-based explanations [17]: example-based explanations [34]: feature importance, SHAP and LIME [35]: feature importance, SHAP [22]: white-box model, insignificant [39]: white-box < black-box, both insignificant [31]: feature importance explanation (transparent system) can be distracting Usability [81]: counterfactuals, pos. for usability [16, 47]: example-based explanations, pos. for satisfaction [67]: CAM-related explanations, pos. for helpfulness [6]: rational-, feature-, example-based explanations, pos. for satisfaction [70]: content-based explanations, pos. for satisfaction [83]: explanations regarding driving information, pos. for ease of use [13]: example-based and rule-based explanations, pos. for helpfulness [71]: local, global, visual (saliency map) explanations, pos. for bug identification [65]: attribution methods and conceptual explanations, pos. for usefulness [84]: feature-based, pos. for reliability [24]: (proposed) template-based expl. pos. for debugging and usefulness [27]: feature importance, counterfactual explanations pos. for perceived fairness [82]: counterfactuals, significant for helpfulness/usability but insignificant for usefulness [1]: ontology-based explanation, insignificant for satisfaction [65]: attribution methods and conceptual explanations, insignificant for ease of use [24]: visual explanations increases usefulness, but improvement is insignificant [3]: pos. for cognitive load/usability (simulation), but insignificant in real-world [29]: feature-based explanations, negative for satisfaction [38]: informing users about the algorithmic decisions, negative ranking scores of recommendations, insignificant for perceived fairness [27]: highlight features only, insignificant for perceived fairness [78]: insignificant in between-subjects but significant in within-subjects for perceived fairness Human-AI Collaboration Performance [88]: textual explanations with domain knowledge (in chess) [90, 25]: feature-based explanations [91]: exampled-based for experts, feature-based for novices [93]: contrastive explanations [13]: example-based and rule-based explanations [95, 96]: example-based explanations, attributions (AI correctness prediction) [97]: important parts in images as explanations [25]: exampled-based, insignificant [89, 15]: feature-based explanations, insignificant
@@
[…]
! Type Paper Detection Result Wrong decisions (failures) [71, 53, 24] High detection rate in [24]; Moderate detection rate (50%) in [71]; Lower detection rate in  [53] Biases in features used by models [57, 71] Moderate detection rate (50%) in [71]; Moderately high detection rate (>50%) Discrimination/Biases in decisions [72] Humans perform well in bias detection (accuracy=88.9%) and bias description (66.7%) Unfairness (perceived) in models [27, 38, 2, 74] [78, 79, 73] Succeed to judge [78, 27, 74]; Not succeed to judge [2]; Not always (no consensus)  [38, 73]

### v5

@@ 1 Introduction
Trust
[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]
[16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]
Understanding subjective [32, 33, 34, 35, 22, 36, 37, 12, 13, 17, 28, 16, 38, 39, 7, 40, 41, 42, 43, 14, 44, 31]
objective [45, 40, 46, 32, 47, 48, 49, 50, 51, 52, 13, 12, 53, 54, 55, 56, 35, 39, 57, 22, 58, 59, 43, 60, 44]
explanation model [61, 62, 63, 64, 46, 21, 65, 49]
Usability workload [48, 21, 16, 3, 66]
helpfulness [46, 48, 56, 65, 67, 13, 68, 45]
satisfaction [47, 16, 7, 18, 6, 69, 1, 29, 19, 70]
undesired behavior detection [53, 24, 71, 57, 72, 38, 73, 74, 75, 27, 76, 2, 77, 78, 79]
ease of use and others [32, 71, 21, 13, 48, 80, 37, 81, 30, 65, 24, 3, 82, 83, 1, 66, 84, 85, 20, 86, 87]
Human-AI Collaboration Performance [88, 10, 89, 25, 90, 91, 92, 13, 15, 93, 56, 30, 94, 39, 25, 53, 29, 95, 96, 43, 97]
@@
Explainable AI User Study
Keywords
XAI, explainable AI,
explanation, explainable,
explanatory, interpretable,
intelligible, black-box,
machine learning,
explainability, interpretability,
intelligibility, explain
attribution, feature
user study, participant ,
human subject,
empirical study,
lab study,
user evaluation,
human evaluation
@@ 3.2 Foundations of User Studies
Based on a data-driven bibliometric analysis of the references in core papers, we highlight significant research topics within the "Foundational Domain" in Figure 1. It is evident that model explanations and interpretability are pivotal components. This includes papers that introduce explanation methods such as LIME [147], SHAP [148], and other attribution methods. These are a frequent subject of study in works measuring understanding and usability. Additionally, convolutional networks, which are commonly employed in experiments, use tools like GradCAM [149] and various saliency maps to generate model explanations. Notably, many research papers appear within the domain of recommender systems, because many XAI user studies are conducted in the context of recommendation solutions. The EU’s General Data Protection Regulation (GDPR) [150] is frequently mentioned in core papers due to the ongoing debate on the right to explanation” [151]. This debate has significantly influenced the shift in modern AI systems towards explainability. While the ultimate consumers of model explanations are humans, well-established research domains that focus on human understanding are underrepresented. For instance, only a few papers related to “Cognition” are cited compared to those on other algorithmic topics. Millecamp et al. [18] suggest enhancing XAI theory with insights from social sciences, including cognitive science and psychology. Given the scant references to psychology, it appears that only a handful of XAI user studies delve into evaluating XAI from a psychological standpoint. We highlight a nascent research domain of XAI frameworks based on human cognition and behavior theories [142]. This theoretical guidance can also offer conceptual tools for better evaluating XAI from user perspectives. More details about common references can be found in Appendix A.1.
@@ 3.3 Impact of User Studies
Figure 1 presents applications that make use (and thus are the consumers) of the findings from core papers. We noticed that studies on user understanding and trust span a wide range of applications. For example, trust is frequently addressed in the contexts of medical diagnosis and transportation, indicating its significance in high-risk scenarios. Recommendation systems emerge as a primary focus in follow-up works. Papers on usability have a significant impact on fields like data visualization, software development, and education. In these areas, models frequently serve as tools to ease the burden on end users. Human-AI collaboration measures particularly promote the further development of robotics and or natural language processing. The prominence of recommendation systems in both foundational works and their impact implies that XAI is an integral component of contemporary recommendation systems. A comprehensive overview of the fundamental works and application domains can be found in Appendix Figure 4.
@@ 4.1 Models and Explanations
White-box Black-box Other
Feature- based local
[21, 154, 48]
[22, 12, 39]
[6, 50]
[21, 49, 55, 45]
[34, 29, 92, 72]
[39, 35, 40]
[47, 65, 42]
[54, 58, 57]
[71, 56, 50]
[40, 41, 89]
[90, 25, 59]
[95, 43, 60]
global
[12, 53, 74]
[21, 50]
[50]
Example-based
[12, 21, 43]
[74, 6, 96]
[52, 17, 57]
[13, 40, 25]
[32]
(generative
models)
Counterfactual
[12, 37]
[82, 21]
[27, 100]
[57, 65]
Concept-based
[61, 62, 71]
[67, 63, 64]
[57, 100]
Other
[11, 88]
[7, 10]
[155, 1, 9, 15]
[3, 13, 51]
[3, 56, 58]
[49, 55, 36]
[28, 85, 16]
[68, 33, 38]*
[76, 8, 23] *
[69, 70]*
[2]
[18, 20, 19] $\dagger$
[26, 84, 20] $\dagger$
[83, 66, 81] $\dagger$
[85, 30, 14] $\dagger$
[91, 5] $\dagger$
TABLE III: Models and explanations in core papers. Papers are categorized according to types of explanations (column) and types of models (row). * denotes papers using recommendation systems as models; $\dagger$ denotes papers proposing novel interpretable interfaces as studied models.
Besides these four main types of explanations, there are other explanations such as rules [11, 88] or game strategies [7, 10] when AI plays games. More details about concrete models and explanations can be found in Appendix B.
@@ 4.2.1 Trust
Trust measures used in much of the existing research can be divided into two groups: self-reported and observed trust [156]. Self-reported trust is commonly measured by asking users to fill out questionnaires whereas observed trust is quantified by humans’ agreement with the model’s decisions. In Table IX in Appendix, trust measures in these two groups are listed. The agreement rate of users with the model decisions is commonly used [9, 12, 11, 25] as a measure of observed trust. Parallel to observed trust measurement, van der Waa et al. [157] ascribe the user’s alignment behaviors to the persuasive power of model explanations, i.e., the capacity to convince users to follow model decisions despite the correctness. As an extension, trust calibration is defined based on this measure. For example, a high agreement rate to wrongly made decisions represents overtrust, while a low agreement rate to correct decisions means undertrust [12]. In self-reported measurements, researchers either utilize well-developed questionnaires or self-designed ones, with the exception of [4] which conducts a semi-structured interview to explore user opinions. Several works [6, 13, 27, 11, 16, 17, 18, 19, 24] propose their own questionnaires. Among these, a subgroup [16, 13, 18, 19, 24] simply asks users to rate a single statement such as “I trust the system’s recommendation/decision”, which is named as one-dimensional trust by [8]. When deploying previously proposed questionnaires [2, 3, 21, 7, 23, 5, 8, 10, 22, 158], Trust in Automation [159] is the most commonly used one, in which the underlying constructs of trust between human and computerized systems are explored.
@@ 4.2.2 Understanding
An important goal of explanation techniques is to foster users’ understanding of complex ML systems. An important separation has to be made between users’ perceived understanding and their actual comprehension of the underlying model, as the two often do not agree [35, 40]. Cheng et al. [22] explicitly differentiate between objective understanding and self-reported understanding, which we term subjective understanding in this work. While subjective understanding is usually measured through questionnaires, measuring objective understanding requires a proxy task where the users’ understanding is put to a test. Additionally, user studies can be run to assess how well users can understand the explanation itself (and not the underlying model). This can be an important sanity check and is particularly used in the domain of conceptual explanations [160, 62], where the intelligibility of concepts needs to be verified. We refer to the third category as understanding of explanations but defer its detailed findings to Appendix C.3.
Objective Understanding. Works in the subdomain of objective understanding deploy proxy tasks to verify users’ understanding of a model’s inner workings. The most commonly considered domain in works on understanding is finance [53, 48, 49, 40, 47, 35, 39] followed by image classification [13, 21, 52]. One of the most critical design choices when assessing objective understanding is the selection of a suitable proxy task. Doshi-Velez and Kim [120] argue that the task should “maintain the essence of the target application” that is anticipated. One of the most prominent tasks is forward simulation [141, 120]. This task demands subjects that are given an input to simulate, i.e., predict, the model’s output. The extent to which participants can successfully provide the model’s output is also referred to as simulatability [141]. However, scholars have designed many more tasks to quantify understanding and applied them across a variety of data modalities (cf. Table VIII in Appendix for an exhaustive listing).
@@ 4.2.3 Usability
Using model explanations to audit models is one purpose of explainability [130]. Some of the surveyed works study how model explanations can assist users in detecting undesired behaviors of models. These issues mainly include (perceived) unfairness in the model decision-making [38, 74, 78, 79], biases in models [72] or features [57], and wrong decisions (failures) [24] in the studied papers. A detailed summary of types of undesired behaviors is listed in Table VI. In the undesired behavior detection, the effectiveness of explanations is evaluated by objective performance measures, such as the number of bugs identified [71], the share of participants that identify a certain bias [57, First Experiment] or by the deviations between model predictions and human predictions for unusual samples [53]. The perception of users regarding fair treatment by a system has primarily been researched in high-stakes applications such as granting loans [27] or granting bail for criminal offenders [74, 75, 73]. For example, [74, 75, 73] investigate the fairness of COMPAS, a commercial criminal risk estimation tool that was used in the US to help make judicial bail decisions. It is also considered in everyday use-cases such as news [38] and music [77] recommendations, or possible career suggestions [76], where a bias in the underlying system can be to the detriment of the user. As the assessment of fairness is a very subjective matter, questions regarding perceived fairness are prevalent, e.g., “how the software made the prediction was fair” [74], which can be answered on 5- or 7-point Likert scales [74, 75, 27, 38, 2, 73]. Among these works, an effective explanation is the one that can either increase or decrease the fairness perceptions, since the aim of explanations is to show fairness or unfairness. An exhaustive overview of measures for usability is given in Table X of the Appendix.
@@ 4.3 Experimental Design and Analysis
Experimental Design
Between-Subjects Within-Subjects Mixed
Papers
[5, 7, 8, 12, 27, 59, 15]
[17, 21, 22, 23, 25, 72]
[28, 11, 40, 46, 32, 95]
[53, 47, 49, 50, 51, 84]
[56, 39, 38, 37, 36, 43]
[90, 54, 30, 29, 92, 96]
[57, 58, 12, 73, 38]
[75, 76, 77, 78, 82]
[1, 3, 4, 19, 9]
[18, 70, 10, 24, 21]
[26, 13, 52, 35, 91]
[57, 78, 71, 81, 93]
[69, 6, 67, 63, 62]
[68, 41, 60, 14, 45]
[2, 66, 13, 52, 16]
[74, 28, 34, 64, 10]
[33, 147, 83, 65, 12]
[61]
@@ 5 Findings of User studies
In this section, we summarize the primary findings from the core papers. Table V lists findings with respect to four measured quantities. To build an overview of the findings, we divide papers according to their evaluation dimensions, i.e., the independent variables in the user studies. When using the presence of explanations as the evaluation aspect, the findings are summarized in Table V. The listed impacts using explanations are to be seen in comparison with a control group without explanations. Effects are divided into two groups: (1) Positive effects, for example, increasing user trust or understanding; (2) Non-positive effects: the effect can be negative, or not significantly positive (neural), or a mixture of different effects (e.g., feature-based explanations have positive effects but counterfactual explanations do not). Beyond the explanations themselves, other possible evaluation dimensions such as that might have an impact on the perception of XAI, for instance, AI technology literacy, model performance, or the dimensionality of the data. Instead of using the mere presence of explanations, many works compare different explanation techniques with each other (see Appendix D for more details).
@@
As various research questions and findings are addressed in 97 core papers, many papers compare explanation types in order to choose a preferable one, it is not possible to cover all results in one table. Based on them, we outline some interesting trends in the effectiveness of explanations on user experience: (1) Explanations are effective in improving users’ subjective understanding; (2) The effectiveness of explanations in increasing user trust and usability of models is not clear; (3) Explanations are not good at convincing users that models are fair; (4) Interactivity of the model has positive impact on user trust, understanding and model usability. The first three statements can validated through the number of papers obtaining positive or non-positive effects in each category, while the last finding is extracted from Table XI in the Appendix, which details findings with on other independent variables. We encourage the reader to consider the short summary of primary findings in the tables and check for further details according to their specific interests. In the following section, we highlight some findings for each category of measurement.
Evaluation Dimension: Explanations
Effect of explanations compared to no explanations
Positive Non-positive / Mixed
Trust
[13]: example-based, rule-based explanations
[16]: example-based explanations for recommendations
[27]: feature importance
[10]: decision-tree explanation for policy
[28]: explanation corpus given by researchers
[25]: feature-based (saliency map), example-based explanations
[8]: explanations for recommendations
[6]: rationale-based, example-based and feature-based (best)
explanations for online symptom checkers
[15]: confidence scores
[3]: positive in simulation
but no improvement in real-word
[1]: explanations for medical suggestions (Doctor XAI [155])
pos.for observed trust but insignificant for reported trust
[12]: feature-based explanations increase appropriate trust
slightly but counterfactual explanations inconclusively
[22, 21, 24]: feature-based explanation, insignificant
[11]: rule-based explanation, insignificant
[15]: Shapley values, insignificant
[29]: feature-based explanation, negative
Understanding Obj.
[53, 22] white-box model
[40] feature importance, LIME (tabular)
[46] counterfactuals+cues (audio)
[50] manipulatability improved by white-box log. reg.
[54] saliency maps (image)
[59] saliency maps for bias detection and strategy identification
[12] counterfactuals+feature importance
[39]: SHAP, negative for black-box model (education domain)
[39]: Insignificant difference btw. black-box and white-box models
[40]: Prototypes, Anchors, LIME on textual data insignificant
[46]: Counterfactuals and Concepts insignificant (audio data)
[50]: Simulatability results insignificant for LIME,
IG, surrogate model on BERT and Logistic Regression Model,
Manipulatability insignificant for BERT
[58]: Insignificant results for GRAD-CAM,
Saliency Map, uncertainty scores in VQA
[59] saliency maps for failure prediction (image)
[60]: saliency maps, negative for a mix of three interpretation techniques in simulation task
Sub.
[13]: example-based, rule-based explanations
[28]: explanation corpus given by researchers
[12]: feature-, example- and counterfactual-based
[38]: explanations provided by [167] for Facebook News Feed
[16]: example- and feature-based explanations
[17]: example-based explanations
[34]: feature importance, SHAP and LIME
[35]: feature importance, SHAP
[22]: white-box model, insignificant
[39]: white-box < black-box, both insignificant
[31]: feature importance explanation (transparent system) can be distracting
Usability
[81]: counterfactuals, pos. for usability
[16, 47]: example-based explanations, pos. for satisfaction
[67]: CAM-related explanations, pos. for helpfulness
[6]: rational-, feature-, example-based explanations,
pos. for satisfaction
[70]: content-based explanations, pos. for satisfaction
[83]: explanations regarding driving information,
pos. for ease of use
[13]: example-based and rule-based explanations,
pos. for helpfulness
[71]: local, global, visual (saliency map) explanations,
pos. for bug identification
[65]: attribution methods and conceptual explanations,
pos. for usefulness
[84]: feature-based, pos. for reliability
[24]: (proposed) template-based expl.
pos. for debugging and usefulness
[27]: feature importance, counterfactual explanations
pos. for perceived fairness
[82]: counterfactuals, significant for helpfulness/usability
but insignificant for usefulness
[1]: ontology-based explanation, insignificant for satisfaction
[65]: attribution methods and conceptual explanations,
insignificant for ease of use
[24]: visual explanations increases usefulness,
but improvement is insignificant
[3]: pos. for cognitive load/usability (simulation),
but insignificant in real-world
[29]: feature-based explanations, negative for satisfaction
[38]: informing users about the algorithmic decisions, negative
ranking scores of recommendations, insignificant for perceived fairness
[27]: highlight features only, insignificant for perceived fairness
[78]: insignificant in between-subjects
but significant in within-subjects for perceived fairness
Human-AI
Collaboration Performance
[88]: textual explanations with domain knowledge (in chess)
[90, 25]: feature-based explanations
[91]: exampled-based for experts, feature-based for novices
[93]: contrastive explanations
[13]: example-based and rule-based explanations
[95, 96]: example-based explanations, attributions (AI correctness prediction)
[97]: important parts in images as explanations
[25]: exampled-based, insignificant
[89, 15]: feature-based explanations, insignificant
@@
Type Paper Detection Result
Wrong decisions
[71, 53, 24]
High detection rate in [24];
Moderate detection rate (50%) in [71];
Lower detection rate in  [53]
Biases in features
used by models
[57, 71]
Moderate detection rate (50%) in [71];
Moderately high detection rate (>50%)
Discrimination/Biases
in decisions
[72]
Humans perform well in
bias detection (accuracy=88.9%) and
bias description (66.7%)
Unfairness (perceived)
in models
[27, 38, 2, 74]
[78, 79, 73]
Succeed to judge [78, 27, 74];
Not succeed to judge [2];
Not always (no consensus)  [38, 73]